OpenStudy (anonymous):

Graph the equation x/2-y/7=1

OpenStudy (anonymous):

y = mx+b put it in this form

OpenStudy (anonymous):

ok but i am trying to remember how to do that lets see here

OpenStudy (anonymous):

what is the lcd?

OpenStudy (anonymous):

i would say 14

OpenStudy (anonymous):

ok multiply everything by it

OpenStudy (anonymous):

2*14=28 and 7*14=98

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

you said multiply

OpenStudy (anonymous):

(x/2)*14=?

OpenStudy (anonymous):

14/2

OpenStudy (anonymous):

how can you simplify it and where did x go

OpenStudy (anonymous):

well i tell you i am not sure. simplify 14/2= 7/2

OpenStudy (anonymous):

i guess